OpenStudy (jemson):

If G is the geometric mean of X and Y, then 1/(G^2 – X^2) + 1/(G^2 – Y^2) =?

OpenStudy (kirbykirby):

Geometric mean of X and Y is \(\sqrt{XY}\), so: \[\frac{1}{G^2-X^2}+\frac{1}{G^2-Y^2}=\frac{1}{\left(\sqrt{XY}\right)^2-X^2}+\frac{1}{\left(\sqrt{XY}\right)^2-Y^2}\\ =\frac{1}{XY-X^2}+\frac{1}{XY-Y^2}\\ =\frac{1}{X(Y-X)}+\frac{1}{Y(X-Y)}\]

OpenStudy (dumbcow):

\[= \frac{Y-X}{XY(Y-X)} \] \[= \frac{1}{XY}\] \[=\frac{1}{G^2}\]